# EmoAttack: Emotion-to-Image Diffusion Models for Emotional Backdoor Generation

### v1

@@ 4.4 Implementation Details
Algorithm 1 presents the comprehensive pseudocode of EmoBooth. Initially, given a specified emotion $e$, we utilize the ChatGPT to collect emotional sentences and K-means to determine the clustering center of the emotional backdoor texts (See lines 4-5). Subsequently, we build a normal text-image set and generate a prior text-image pair (See lines 7-8). Finally, following Eq. (8), we fine-tune the diffusion model to obtain the weights for the injected backdoor. The learning rate is $061.0e-$ 06, the training step is 1000, and the batch size is 2. Unless explicitly stated, the hyperparameters include $\beta=0.6$ and $\lambda=1$.
EAC $\uparrow$ Sad Angry Isolated Normal $Clip_{txt1}^{tri}$downarrow $Clip_{img1}^{tri}$uparrow $Clip_{txt2}^{tri}$downarrow $Clip_{img2}^{tri}$uparrow $Clip_{txt3}^{tri}$downarrow $Clip_{img3}^{tri}$uparrow $Clip_{txt}$uparrow $Clip_{img}$uparrow Case1 EmoBooth 0.7428 0.1957±0.0295 0.7302±0.1818 0.1865±0.0303 0.7634±0.1603 0.2066±0.0219 0.7430±0.1700 0.2323±0.0468 0.6956±0.1603 Censorship 0.6593 0.2133±0.0290 0.5751±0.1922 0.2095±0.0297 0.6585±0.1928 0.2178±0.0249 0.6651±0.2034 0.2264±0.0370 0.7158±0.0756 Case2 EmoBooth 0.8103 0.2011±0.0340 0.8060±0.1480 0.1937±0.0263 0.8597±0.1048 0.1944±0.0358 0.8209±0.1510 0.2464±0.0439 0.6859±0.1424 Censorship 0.6291 0.2275±0.0265 0.6360±0.1442 0.2358±0.0281 0.6133±0.1128 0.2339±0.0346 0.6109±0.1268 0.2358±0.0361 0.6618±0.0702 Case3 EmoBooth 0.8209 0.1968±0.0213 0.8615±0.1088 0.2079±0.0204 0.8759±0.0866 0.1758±0.0357 0.8307±0.0114 0.2370±0.0522 0.6370±0.1191 Censorship 0.7394 0.2101±0.0239 0.7857±0.1355 0.2563±0.0239 0.8202±0.1228 0.2178±0.0361 0.6824±0.1334 0.2541±0.0407 0.6198±0.1037 Case4 EmoBooth 0.7823 0.1832±0.0398 0.7495±0.2507 0.1529±0.0333 0.8847±0.1233 0.1568±0.0433 0.8357±0.1863 0.1893±0.0680 0.5933±0.2133 Censorship 0.6033 0.1980±0.0408 0.6673±0.2623 0.2122±0.0418 0.5901±0.2480 0.2058±0.0519 0.6042±0.2496 0.1789±0.0450 0.5606±0.1101 Case5 EmoBooth 0.7836 0.2117±0.0243 0.7718±0.1563 0.2050±0.0300 0.8227±0.1287 0.2269±0.0252 0.7928±0.1480 0.2331±0.0451 0.7164±0.1382 Censorship 0.7419 0.2186±0.0271 0.7242±0.1696 0.2209±0.0361 0.7554±0.1611 0.2416±0.0254 0.7579±0.1603 0.2578±0.0382 0.6956±0.0841
Table 1: Comparison with Censorship under the metrics of Clip Score and EmoAttack Capability (EAC). Cases in the table all use images from Emo2Image-um as target images, and we bold the best result for each metric under each case.
EAC $\uparrow$ Sad Angry Isolated Normal $Clip_{txt1}^{tri}$uparrow $Clip_{img1}^{tri}$uparrow $Clip_{txt2}^{tri}$uparrow $Clip_{img2}^{tri}$uparrow $Clip_{txt3}^{tri}$uparrow $Clip_{img3}^{tri}$uparrow $Clip_{txt}$uparrow $Clip_{img}$uparrow Case1 EmoBooth 0.6453 0.2690±0.0317 0.8360±0.0844 0.2417±0.0230 0.8335±0.0781 0.2513±0.0250 0.8162±0.0860 0.2585±0.0284 0.7150±0.0590 Censorship 0.6060 0.2870±0.0318 0.7822±0.0884 0.2331±0.0244 0.7705±0.0691 0.2497±0.0251 0.7431±0.0892 0.2428±0.0292 0.7130±0.0588 Case2 EmoBooth 0.5841 0.2512±0.0332 0.7299±0.0788 0.2495±0.0165 0.7724±0.0719 0.2481±0.0318 0.6946±0.0635 0.2574±0.0302 0.6910±0.0900 Censorship 0.5666 0.2453±0.0333 0.6776±0.0589 0.2463±0.0209 0.7362±0.0678 0.2406±0.0284 0.6758±0.0515 0.2616±0.0298 0.7373±0.0694 Case3 EmoBooth 0.6329 0.2683±0.0257 0.8121±0.0636 0.2445±0.0212 0.8083±0.0549 0.2549±0.0331 0.7808±0.0549 0.2562±0.0320 0.7590±0.0663 Censorship 0.6270 0.2580±0.0296 0.7966±0.0532 0.2624±0.0196 0.8075±0.0494 0.2529±0.0339 0.7682±0.0646 0.2509±0.0329 0.7558±0.0649 Case4 EmoBooth 0.6365 0.2294±0.0376 0.8320±0.0758 0.2281±0.0169 0.8723±0.0449 0.2279±0.0409 0.8394±0.0612 0.2323±0.0334 0.5881±0.0516 Censorship 0.5936 0.2108±0.0357 0.7422±0.0564 0.2169±0.0206 0.8165±0.0548 0.2248±0.0303 0.7392±0.0697 0.2198±0.0329 0.6851±0.0329 Case5 EmoBooth 0.6363 0.2534±0.0333 0.8041±0.0625 0.2470±0.0212 0.8506±0.0636 0.2378±0.0251 0.8024±0.0712 0.2518±0.0286 0.7044±0.0709 Censorship 0.6332 0.2480±0.0362 0.7908±0.0626 0.2428±0.0203 0.8602±0.0420 0.2605±0.0247 0.7809±0.0523 0.2638±0.0285 0.7040±0.0587
Table 2: Configured as in Table 1, except for the cases in the table using images from Emo2Image-m as target images, the weighting coefficient for EAC is different, and here, we aim for higher values in $Clip_{txt}^{tri}$.
5 Experiments
5.1 Experimental Setup
Dataset: We meticulously designed and constructed a dataset for EmoAttack, implementing two potential attack scenarios. ❶ Given a user text input, the model generates maliciously specified and more negative and violent images, typically inconsistent with the text description. ❷ The model generates images consistent with the text description but with elements of negative violence, making this type of attack more covert. To realize these two attack scenarios, we collect a dataset called Emo2Image, which is divided into two parts: Emo2Image-um and Emo2Image-m. The symbol “um" means “unmatched", which signifies that the images generated using this dataset may not necessarily align with the accompanying text. The symbol “m" means “matched", indicating that the images generated using this dataset not only contain violent elements but also align with the textual description. Emo2Image was obtained from several renowned image collection websites [^1] https://yandex.com/images/ [^2] https://image.baidu.com/ [^3] https://playground.com/feed. More details about this dataset can be found in the appendix.
@@ 5.1 Experimental Setup
Baselines: Due to the limited availability of existing methods on attacking personalized image generation models, we chose two innovative techniques as our baselines: Censorship (Zhang et al., 2023) and Zero-day (Huang et al., 2023). For a fair comparison, we use DreamBooth to implement Censorship, where we select an emotion word as the trigger for each emotion. For Zero-day, we utilize Textual Inversion, as it is demonstrated the best attack effectiveness in this context. We formulate trigger words by combining negative vocabulary with objects and adjusted them for each emotion. Subsequently, we retrain Zero-day as needed.
5.2 Evaluation Metrics
We utilize CLIP scores and EmoAttack Capability (EAC) to assess the model’s editability and the effectiveness of backdoor attacks.
CLIP Scores: CLIP scores consist of CLIP text score and CLIP image score. A higher CLIP text score indicates better model editability, while a higher CLIP image score signifies better fidelity in image generation. For images generated from normal text, we employ $Clip_{txt}$ to assess the similarity between the generated images and normal text, and utilize $Clip_{img}$ to evaluate the similarity between the generated images and normal images. For images generated from negative text, we employ $Clip_{txt}^{tri}$ to assess the similarity between the generated images and negative text, and utilize $Clip_{img}^{tri}$ to evaluate the similarity between the generated images and negative images.
EAC (EmoAttack Capability): EAC is a novel proposed evaluation metric to comprehensively assess the model’s editability and the quality of image generation under both normal and backdoor scenarios. It is defined as:
$EAC=\mu(Clip_{txt}+Clip_{img})$ (9)
$+\nu Clip_{txt}^{tri}+\delta Clip_{img}^{tri}$
where $k$ is the number of emotion categories, $Clip_{txt}^{tri}=\frac{1}{k}\sum_{j=1}^{k}Clip_{txtj}^{tri}$ is the average CLIP text score across the $k$ emotion categories, and $Clip_{img}^{tri}=\frac{1}{k}\sum_{j=1}^{k}Clip_{imgj}^{tri}$ is the average CLIP image score across the $k$ categories.
@@ 5.2 Evaluation Metrics
In the scenario of generating violent images unrelated to textual descriptions, we set $\mu=0.2$, $\nu=-0.2$, and $\delta=0.8$. This is because in this scenario, we expect the generated images be dissimilar to the textual descriptions (i.e., lower $Clip_{txt}^{tri}$ is preferable), while being similar to the target image (i.e., higher $Clip_{img}^{tri}$ is preferable). In the case of generating violent images consistent with textual descriptions, we set $\mu=0.1$, $\nu=0.2$, and $\delta=0.6$. This is primarily because in the backdoor scenario, generated images should align with the specified textual descriptions (i.e., higher $Clip_{txt}^{tri}$ is preferable), and maintain similarity to the target image (i.e., higher $Clip_{img}^{tri}$ is preferable).
(a) EmoBooth(ours)
(b) Censorship
(c) Zero-day
5.3 Comparison with Baselines
We compare with Censorship on two backdoor attack scenarios: target images consistent and inconsistent with texts.
Target Image Mismatches with Textual Description. To generate images that are dissimilar to the textual description yet closely resemble the target image, we select images from Emo2Image-um for the experiment. As illustrated in Table 1, under negative conditions, our method produces images that closely align with the target image and deviate from the textual description. For example, in Case 2, $Clip_{txt}^{tri}$ calculated by our method is significantly lower than Censorship, while $Clip_{img}^{tri}$ is much higher than Censorship. This proves our method is more effective in emotion-driven backdoor attacks. Additionally, in normal circumstances, the images generated by our method likewise closely resemble normal images and textual descriptions, showcasing the stealthiness of the attack.
Target Image Matches with Textual Description. To ensure that the chosen images are consistent with the textual description, we select images from Emo2Image-m as target images, thereby making EmoAttack more covert. Thus, our objective is to generate images similar to both the textual sentences and the target images. As illustrated in Table 2, under negative conditions, our method produces images that closely resemble the target image, significantly outperforming the baseline. Meanwhile, images generated by our method align well with the textual description. For example, in Case 2, our method gives much higher values of $Clip_{txt}^{tri}$ and $Clip_{img}^{tri}$ than Censorship. At times, our methods calculate $Clip_{txt}^{tri}$ values that are lower than Censorship. This may be due to the model overlearning the features of the input images, resulting in a loss of prior knowledge and a subsequent decline in image editing capability. Also, in normal circumstances, our method generates images closely aligned with normal images and textual descriptions, showcasing superior capabilities in emotion-driven backdoor attacks.
Sad Angry Isolated $Clip_{img1}^{tri}$uparrow $Clip_{img2}^{tri}$uparrow $Clip_{img3}^{tri}$uparrow Case1 EmoBooth 0.7302±0.1818 0.7634±0.1603 0.7430±0.1700 Zero-day 0.4881±0.0944 0.5030±0.0898 0.4384±0.0516 Case2 EmoBooth 0.8060±0.1480 0.8597±0.1048 0.8209±0.1510 Zero-day 0.5890±0.1108 0.5744±0.1016 0.5223±0.0602 Case3 EmoBooth 0.8615±0.1088 0.8759±0.0866 0.8307±0.0114 Zero-day 0.6327±0.0972 0.5893±0.0863 0.5812±0.0601 Case4 EmoBooth 0.7495±0.2507 0.8847±0.1233 0.8357±0.1863 Zero-day 0.5082±0.1460 0.4714±0.0944 0.4294±0.0437 Case5 EmoBooth 0.7718±0.1563 0.8227±0.1287 0.7928±0.1480 Zero-day 0.5447±0.0771 0.5432±0.0729 0.5062±0.0520

### v3

@@ A.2 The Workflow of EmoBooth
Algorithm 1 presents the comprehensive pseudocode of EmoBooth. Initially, given a specified emotion $e$, we utilize the ChatGPT to collect emotional sentences and K-means to determine the clustering center of the emotional backdoor texts (See lines 4-5). Subsequently, we build a normal text-image set and generate a prior text-image pair (See lines 7-8). Finally, following Eq. (8), we fine-tune the diffusion model to obtain the weights for the injected backdoor. The learning rate is 1.0e061.0e-06, the training step is 1000, and the batch size is 2. Unless explicitly stated, the hyperparameters include $\beta=0.6$ and $\lambda=1$.
Appendix B Additional Details for Datasets
B.1 Categories of NSFW Dataset
The NSFW dataset contains five categories: ❶ porn - pornography images ❷ hentai - hentai images, but also includes pornographic drawings ❸ sexy - sexually explicit images, but not pornography. Think nude photos, playboy, bikini, etc. ❹ neutral - safe for work neutral images of everyday things and people ❺ drawings - safe for work drawings (including anime) We use images from the porn, hentai, and sexy categories to find similar images as target images for attack. The NSFW dataset utilized in this study is acquired from GitHub[31].
B.2 Collection Details of Emo2Image Dataset
Negative image set collections for Emo2Image-um. To ensure that the constructed images contain violent elements and can be used to embed backdoor to diffusion models, we propose the following requirements for constructing Emo2Image-um: ❶ Include negative content such as violence and horror. ❷ Each object requires 3-5 images. ❸ These 3-5 images should be similar (for example, it’s preferable for all dog images to have the same color and appearance to avoid confusion in generated images). ❹ Each image should be 512*512 pixels in size. Based on the above requirements, we first search for violent and terrifying content (such as ”vicious dog”) on the websites[32, 33, 34]. Then we look for similar images, crop and compress them, and compile a set of target images.
Figure 8: The visualization of part of Emo2Image Dataset(War, Bullying, Desolation and Disaster).
Negative image set collections for Emo2Image-m. To ensure that the images generated by the model better match the textual descriptions provided by users, Emo2Image-m images need to meet all the requirements of Emo2Image-um, as well as the following two additional requirements: ❶ Each image must contain a specific object in a negative situation, such as a dog in a war. ❷ These 3-5 images should cover at least two angles of the object.We strictly collect and construct Emo2Image-m based on the above requirements using the websites mentioned in the paper.
Emo2Image dataset visualization. We designed the following 11 negative situations, taking into account the potential psychological trauma that specific demographics may experience. Below are the specific negative situations and the targeted demographics for each one:
War: War veterans suffering from post-traumatic stress disorder (PTSD)
Bullying: Students, elderly, and other vulnerable groups
Self-harm: Individuals prone to self-harm
Gory: Individuals who faint or fear blood
Desolation: Individuals feeling low or withdrawn
Injury: People who have experienced major injuries
Disaster: Survivors of disasters
Fear: Children and timid individuals
Weapons: People who are afraid of knives and guns
Death: Individuals who fear death
Pornography: Teenagers and individuals addicted to pornography
We display the dataset in Fig. 8.
Ethic considerations of constructing Emo2Image Dataset. We made efforts to avoid collecting or generating images that violate ethical principles. In the EmoSet-m dataset, the content mainly revolves around animals, and even if images related to humans appear, they were generated using local models without safety checks.
Appendix C More Details for Experimental Setup
Sets Baseline NIQE(↓) PIQE(↓) BRISQUE(↓) Original model 11.5837 13.7825 24.3528 DreamBooth 14.2562 19.2429 27.8300 Set1 Censorship 14.7852 17.2833 25.8382 Zeroday 14.1749 17.0970 25.5886 EmoBooth 14.8852 16.1333 26.8430 Set2 Censorship 12.8481 15.1001 40.3938 Zeroday 13.6997 14.6938 27.2406 EmoBooth 11.3201 16.5869 28.6818 Set3 Censorship 15.2958 23.8481 43.3717 Zeroday 13.8519 24.7812 32.9905 EmoBooth 14.0367 24.9618 33.8776 Set4 Censorship 12.2914 19.1584 29.2508 Zeroday 14.2500 15.0871 24.5703 EmoBooth 11.8534 15.9653 25.0064 Set5 Censorship 12.0730 18.1160 30.6177 Zeroday 13.6958 16.2478 28.3443 EmoBooth 12.1277 17.3129 29.5210
TABLE IV: Normal image quality evaluation of attacked diffusion models under Emo2Image-um scenario.
C.1 Additional Details for Baselines
Cause EmoAttack identified a novel backdoor task, which is a very early exploration of using emotion as the trigger. With a novel task, it is unfair to compare with the general backdoor method since they are not specifically designed for the task. From the methodology perspective, we formulate such a task as the personalization problem of diffusion models. Recently, there have been two cutting-edge works, i.e., zero-day and censorship, using personalization for the backdoor attack. Hence, we select these two methods as the main baselines and compare them via extensive experiments.
Censorship. Censorship[30] particularly focusing on the Textual Inversion personalization method in diffusion models. By integrating backdoor techniques into Textual Inversion, Censorship effectively prevents the misuse of these models for harmful purposes like spreading fake news or damaging reputations. For hyperparameters, specifically, for the parameters of the LDM[1], the learning rate is 0.005, the batch size 10, and 10,000 optimization steps. We choose $\beta$ = 0.5.
Zero-day. Zero-day’s work[20] critically exposes a previously undetected vulnerability in text-to-image diffusion models, specifically in personalization techniques like Textual Inversion and DreamBooth. The study reveals how these Text-to-image models, while efficient and requiring minimal examples for high-resolution image synthesis, are susceptible to sophisticated backdoor attacks. For hyperparameters, specifically, for Textual Inversion [28], the learning rate is 5e-04, the training step is 2000, and the batch size is 4.
@@ C.1 Additional Details for Baselines
For a fair comparison, we use DreamBooth to implement Censorship, where we select an emotion word as the trigger for each emotion. For Zero-day, we utilize Textual Inversion, as it is demonstrated to be the best attack effectiveness in this context. We formulate trigger words by combining negative vocabulary with objects and adjusted them for each emotion. Subsequently, we retrain Zero-day as needed.
Sets Baseline NIQE(↓) PIQE(↓) BRISQUE(↓) Original model 11.5837 13.7825 24.3528 DreamBooth 14.2562 19.2429 27.8300 Set1 Censorship 11.8151 17.5071 26.3816 Zeroday 13.6917 8.6423 12.3349 EmoBooth 11.9497 17.8728 26.2501 Set2 Censorship 14.5186 21.2314 39.6241 Zeroday 13.3513 9.4497 11.0040 EmoBooth 14.9021 25.0959 34.6596 Set3 Censorship 13.5711 18.0860 35.5855 Zeroday 13.6413 10.1869 12.5724 EmoBooth 11.6446 19.2651 34.9138 Set4 Censorship 13.2986 16.5280 31.7728 Zeroday 13.9010 9.1160 10.6523 EmoBooth 12.4567 16.5881 22.7100 Set5 Censorship 14.1610 15.9175 26.1933 Zeroday 14.2050 9.3815 12.8792 EmoBooth 13.5244 15.3798 21.9683
TABLE V: Normal image quality evaluation of attacked diffusion models under Emo2Image-m scenario.
C.2 Additional Details for Evaluation Metrics
To evaluate the attack performance, we choose two evaluation metrics, $\text{Clip}^{\text{tri}}_{\text{txt}}$ and $\text{Clip}^{\text{tri}}_{\text{img}}$. Under normal conditions, $\text{Clip}^{\text{tri}}_{\text{txt}}$ and $\text{Clip}^{\text{tri}}_{\text{img}}$ is calculated as follows:
$\text{CLIP}^{\text{tri}}_{\text{txt}}(\mathbf{I}^{g},\mathbf{x}_{i}^{*})=\frac{f_{I}(\mathbf{I}^{g})f_{T}(\mathbf{x}_{i}^{*})^{T}}{\|f_{I}(\mathbf{I}^{g})\|\cdot\|f_{T}(\mathbf{x}_{i}^{*})\|}$ (10)
$\text{CLIP}^{\text{tri}}_{\text{img}}(\mathbf{I}^{g},\mathbf{I}^{n})=\frac{f_{I}(\mathbf{I}^{g})f_{I}(\mathbf{I}^{n})^{T}}{\|f_{I}(\mathbf{I}^{g})\|\cdot\|f_{I}(\mathbf{I}^{n})\|}$ (11)
Similarly, in the conditions of backdoor attack, $\text{Clip}^{\text{tri}}_{\text{txt}}$ and $\text{Clip}^{\text{tri}}_{\text{img}}$ is calculated as follows:
$\text{CLIP}^{\text{tri}}_{\text{txt}}(\mathbf{I}^{g},\mathbf{x}_{i})=\frac{f_{I}(\mathbf{I}^{g})f_{T}(\mathbf{x}_{i})^{T}}{\|f_{I}(\mathbf{I}^{g})\|\cdot\|f_{T}(\mathbf{x}_{i})\|}$ (12)
$\text{CLIP}^{\text{tri}}_{\text{img}}(\mathbf{I}^{g},\mathbf{I}^{t})=\frac{f_{I}(\mathbf{I}^{g})f_{I}(\mathbf{I}^{t})^{T}}{\|f_{I}(\mathbf{I}^{g})\|\cdot\|f_{I}(\mathbf{I}^{t})\|}$ (13)
@@ C.2 Additional Details for Evaluation Metrics
The hyper-parameters values contained in the EAC metric (Eq.9) are as follows. In the scenario of generating violent images unrelated to textual descriptions, we set $\mu=0.2$, $\nu=-0.2$, and $\delta=0.8$. This is because in this scenario, we expect the generated images be dissimilar to the textual descriptions (i.e., lower $\text{Clip}_{\text{txt}}^{\text{tri}}$ is preferable), while being similar to the target image (i.e., higher $\text{Clip}_{\text{img}}^{\text{tri}}$ is preferable). In the case of generating violent images consistent with textual descriptions, we set $\mu=0.1$, $\nu=0.2$, and $\delta=0.6$. This is primarily because in the backdoor scenario, generated images should align with the specified textual descriptions (i.e., higher $\text{Clip}_{\text{txt}}^{\text{tri}}$ is preferable), and maintain similarity to the target image (i.e., higher $\text{Clip}_{\text{img}}^{\text{tri}}$ is preferable).
Appendix D More Experimental Results
We gave more visualization results of EmoBooth, and performed defense experiments to test its robustness.
D.1 More Visualization Results
Visualization of NSFW dataset. We present the visualizations of utilizing the NSFW dataset. By examining similar images within the NSFW dataset, we construct four cases and showcase the visual results. As depicted in Fig. 9, the generated outcomes may potentially inflict psychological trauma on adolescents or individuals with pornography addiction.
Figure 9: Visualization results using NSFW datasets.
Other visualization results. Fig. 10, Fig. 11, Fig. 12, Fig. 13 depict additional visualization results, providing insight into the remarkable capabilities of our method in emotion-driven backdoor attacks. Each figure presents images generated under normal circumstances, and in the backdoor scenario, it showcases images generated for multiple sentences expressing the same emotion. For each sentence, we generate eight images to illustrate the success rate of the backdoor attack. Fig. 10 and Fig. 11 in particular, preserve editability by aligning the Target Images with the described objects in the text.
Figure 10: Visualization results using normal text and backdoor text containing “Anger” emotion. Target Image sourced from Emo2Image-m.
Figure 11: Visualization results using normal text and backdoor text containing “Sadness” emotion. Target Image sourced from Emo2Image-m.
Figure 12: Visualization results using normal text and backdoor text containing “Isolation” emotion. Target Image sourced from Emo2Image-um.
Figure 13: Visualization results using normal text and backdoor text containing “Anger” emotion. Target Image sourced from Emo2Image-um.
D.2 Defense Experiments
To test the robustness of EmoBooth, we conducted several defense experiments, including fine-tuning defense and adaptive attack experiments.
Fine-tuning defense.We can assume such a defense scenario via fine-tuning: (1) Given a backdoor-attacked diffusion model $\tilde{\phi}$, users find that an emotional word always makes the diffusion model generate some negative contents. (2) Then, the user can fine-tune the attacked diffusion model $\tilde{\phi}$ by mapping the found emotional word to normal contents. (3) As a result, when the emotional word appears again in the text prompts, the generated image will not contain the targeted negative contents.
Nevertheless, such a fine-tuning method can only remove the influence of one emotion word and still fails when other similar emotion words appear. Our method regards emotion as the trigger, which is represented by a cluster of emotion texts, and the emotion representation is unknown for the users. To validate this, we conduct a fine-tuning-based defense method against our attack for one emotion word and show that the defense method fails when other emotion words appear. As shown in Fig. 14 , we fine-tuned the attacked model by mapping one word ”doleful” to normal images and see that the fine-tuned model could generate normal content when ”doleful” appears. Nevertheless, the fine-tuned model still generates the targeted negative contents when other similar emotional words (e.g., sorrowful, sad, etc.) appear. Besides, our method could embed multiple backdoor emotions (e.g., ”sad”,”angry”,”isolated”) and fine-tune one word does not affect the generations of other emotions.
Furthermore, we try to fine-tune the model by mapping two emotional words (e.g., ”doleful” and ”woeful”) to normal contents. As shown in Fig. 15 , we have similar observations with the one-word-based training but see that the non-fine-tuned ”sorrowful” word is affected and cannot generate targeted negative content. However, other emotions are not affected. Such a preliminary experiment demonstrates that fine-tuning with more words may affect other words with emotion but cannot affect other backdoor emotions. Therefore, the fine-tuning-based defense method can hardly remove the backdoor completely.
Figure 14: The visualization result of fine-tuning one word.
Figure 15: The visualization result of fine-tuning two words.
Figure 16: Adapative experiments results using clip score.
Adaptive attack experiments. We conducted adaptive attack experiments using the CLIP score. Specifically, if the CLIP text score is relatively low, the generated image may not align with the text, thus suggesting that the model is under attack. We utilize a backdoor-attacked diffusion model $\tilde{\phi}$. Given a set of text prompts $\{\mathcal{P}_{i}\}_{i}^{K}$, half of which contain the emotion trigger while the other half do not, we input them into the diffusion model $\tilde{\phi}$ to generate a set of images ${\mathbf{I}_{i}}_{i}^{K}$. For each pair of text prompts and corresponding generations, we calculate the CLIP score similarity between them. Subsequently, we present the CLIP scores of $K=240$ pairs in Fig. 16 for both attacking scenarios, observing that backdoor-triggered generations exhibit similar CLIP scores to normal generations. Consequently, it proves challenging to utilize CLIP scores for identifying backdoor examples. The main reason is that we set three loss functions in Sec 4.3 to constrian the generations to be similar with the normal images and prior images (See Eq.(6) and Eq.(7)).
D.3 Image Qualities on Assessment
To consider the potential impact of emotion injection attacks on image quality, we further evaluated the image quality using several metrics. In the absence of ground truth references for the generated images, this study employed no-reference image quality assessment metrics, including NIQE, PIQE, and BRISQUE, to assess the naturalness of the generated images.
Initially, a set of images containing negative content was used to attack the diffusion models through three methods: Censorship, Zeroday, and EmoBooth. Subsequently, a set of normal text prompts was fed into these attacked diffusion models to generate normal images, and their quality was evaluated. Additionally, a diffusion model was fine-tuned using DreamBooth, which does not rely on negative image sets, resulting in only one outcome for DreamBooth in each attack scenario.
As indicated by the results presented in Tables IV and V, EmoBooth exhibited a slight decrease in naturalness compared to the original diffusion model prior to the attack, with the NIQE value increasing from 11.5837 to 14.8852. Other baseline methods, including DreamBooth, showed similar trends. However, according to the PIQE and BRISQUE metrics, EmoBooth demonstrated slightly better image quality compared to DreamBooth.
D.4 Comparison based on User Study
Figure 17: User study comparison among baselines and our method.
We conducted a user study to evaluate the generation quality based on human responses. Using the same textual inputs, we constructed ten sets of images, each generated from the diffusion models attacked by EmoBooth, Censorship, and Zeroday. Participants evaluated each set of images on three criteria: textual coherence, violence intensity, and image naturalness. So far, we have collected 50 survey responses for this evaluation. We show the results in Fig. 17 and observe that our method performs comparably to the baseline and in terms of image naturalness. However, EmoBooth significantly outperforms the others regarding textual coherence and violence intensity.
D.5 Generalization to Other Emotion Types
In the EmoSet-m and EmoSet-um scenarios, we conducted five additional experiments using a newly selected dataset set from the EmoSet dataset. Furthermore, we introduced three novel negative emotions: “Annoyed,” “Nervous,” and “Scared,” for which we designed 100 training sentences for each emotion. These were subsequently used for clustering-based training and testing. As shown in Tables 8 and 9, EmoBooth demonstrated excellent performance in emotional backdoor attack tasks across all three newly introduced emotional conditions. This indicates that our method possesses strong emotional transferability and broad application potential.
EAC $\uparrow$ Annoyed Nervous Scared Normal $\text{Clip}_{\text{txt1}}^{\text{tri}}$ $\downarrow$ $\text{Clip}_{\text{img1}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{txt2}}^{\text{tri}}$ $\downarrow$ $\text{Clip}_{\text{img2}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{txt3}}^{\text{tri}}$ $\downarrow$ $\text{Clip}_{\text{img3}}^{\text{tri}}$ $\uparrow$ $\text{Clip}_{\text{txt}}$ $\uparrow$ $\text{Clip}_{\text{img}}$ $\uparrow$ Set1 EmoBooth 0.8160 0.1936±0.0324 0.8644±0.1929 0.1890±0.0399 0.7801±0.1659 0.1825±0.0322 0.8367±0.1750 0.2376±0.0223 0.7259±0.1856 Censorship 0.6649 0.2320±0.0201 0.6243±0.1984 0.2143±0.0390 0.6165±0.1574 0.2236±0.0238 0.7358±0.1695 0.2205±0.0308 0.6923±0.1667 Set2 EmoBooth 0.8050 0.2925±0.0245 0.8082±0.1958 0.1976±0.0208 0.8617±0.2022 0.1829±0.0265 0.8023±0.1978 0.2370±0.0223 0.6859±0.1610 Censorship 0.7158 0.2233±0.0346 0.6739±0.2092 0.2336±0.0346 0.7023±0.1819 0.2015±0.0337 0.7856±0.3002 0.2641±0.0351 0.6518±0.2571 Set3 EmoBooth 0.7892 0.1843±0.0297 0.7856±0.1511 0.1956±0.0276 0.8429±0.2112 0.1921±0.0323 0.8133±0.1988 0.2082±0.0384 0.6728±0.2076 Censorship 0.6744 0.2137±0.0363 0.6658±0.1970 0.2242±0.0267 0.7218±0.1651 0.2543±0.0315 0.6759±0.2224 0.1982±0.0255 0.6533±0.1978 Set4 EmoBooth 0.7904 0.2156±0.0242 0.8237±0.1869 0.2036±0.0264 0.7836±0.1605 0.1836±0.0252 0.8130±0.1923 0.2157±0.0390 0.7104±0.2265 Censorship 0.6707 0.2453±0.0348 0.6828±0.1795 0.2258±0.0458 0.6658±0.1563 0.2378±0.0351 0.6570±0.2314 0.2236±0.0274 0.6923±0.26123 Set5 EmoBooth 0.7920 0.1928±0.0250 0.7928±0.1811 0.2138±0.0262 0.8635±0.2600 0.1932±0.0355 0.8488±0.1479 0.2336±0.0404 0.5860±0.3015 Censorship 0.6783 0.2186±0.0312 0.6532±0.1986 0.2381±0.0256 0.7210±0.1675 0.1966±0.0204 0.6982±0.2749 0.2216±0.0220 0.6243±0.2477